OpenStudy (anonymous):

What is the solution to the system of equations? 5x-3y=-9 2x-5y=4 will fan and medal

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

x=-3 y=-2

OpenStudy (igreen):

@Aleah.wood0613 Please do not give answers. Read the code-of-conduct: http://openstudy.com/code-of-conduct

OpenStudy (solomonzelman):

my favorite way since algebra 2 class. \(\large\color{black}{ \left[\begin{matrix}5 & ~~-3&~~-9 \\ 2 & ~~-5&~~4\end{matrix}\right] }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}5 & ~~-3&~~~~~-9 \\ 2 & ~~-5&~~~~~~4\end{matrix}\right] }\) \(\large\color{black}{ \left[\begin{matrix}10 & ~~-6&~~~~~-18 \\ -10 & ~~25&~~~~~~-20\end{matrix}\right] }\) \(\large\color{black}{ \left[\begin{matrix}10 & ~~-6&~~~~~-18 \\ 0 & ~~19&~~~~~~-38\end{matrix}\right] }\) \(\large\color{black}{ \left[\begin{matrix}10 & ~~-6&~~~~~-18 \\ 0 & ~~1&~~~~~~-2\end{matrix}\right] }\)

OpenStudy (solomonzelman):

now you need the x....

OpenStudy (solomonzelman):

this is what we do with the first row: \(\large\color{black}{ \left[\begin{matrix} 0 & ~~1&~~~~~~-2\end{matrix}\right] }\) times 6, >> \(\large\color{black}{ \left[\begin{matrix} 0 & ~~6&~~~~~~-12\end{matrix}\right] }\) add this to the 1st row: \(\large\color{black}{~~~ \left[\begin{matrix} 10 & ~~-6&~~~~~-18\end{matrix}\right] }\) \(\large\color{black}{ ^{^+}\left[\begin{matrix} 0 & ~~~~~~~~~6&~~~~~~-12\end{matrix}\right] }\) ------------------------------- \(\large\color{black}{ \left[\begin{matrix} 0 & ~~0&~~~~~~-30\end{matrix}\right] }\)

OpenStudy (solomonzelman):

I meant 10, 0 -30 for the very last thing.

OpenStudy (solomonzelman):

So we get; \(\large\color{black}{ \left[\begin{matrix}10 & ~~0&~~~~~-30\\ 0 & ~~1&~~~~~~-2\end{matrix}\right] }\) dividing the 1st row by 3, \(\large\color{black}{ \left[\begin{matrix}1 & ~~0&~~~~~-3\\ 0 & ~~1&~~~~~~-2\end{matrix}\right] }\)

OpenStudy (solomonzelman):

I mean once the answer was given away I went ahead and verified it

OpenStudy (anonymous):

thanks for explaining it better

OpenStudy (solomonzelman):

yw, if that expln. is a satisfactory